if you invested $30 in a single stock 40 years ago and it grew 25% annually, what would the stock be worth today? use the exponential growth equation
Answer #1

Exponential growth formula

Y = I * (1 + R)^T

Y = Stock worth Today

I = Intial Value

R = Growth Rate

T= Time Period

Y = 30(1+ 0.25)^40 = 2,25,694.915

ALL THE BEST :)))

In case of any query feel free to leave your doubt in comment section.